# The dawn is quiet II: Gaia XP constraints on the Milky Way’s proto-Galaxy from very metal-poor MDF tails

### v1

@@ 4.1 Zero Rotation Expected for the Proto-Galaxy
Figure 7: Median MDFs from one-zone GCE models grouped by exponential density slope $k=0.6,1.0,1.4$. The number of models within each slope category is annotated. Lines and shaded regions show the median and 16th–84th percentile ranges, respectively.
@@
Figure 10: The cumulative stellar mass assembly history as a function of metallicity ([Fe/H]) for three slope classes in our GCE models. Lines and shaded regions show the median and 16th–84th percentile ranges, respectively. 2024arXiv241022250K estimated the mass of the in situ proto-Galaxy to be between $1.07\times 10^{7}$ and $2.11\times 10^{7}\,\mathrm{M_{\odot}}$ between [Fe/H] = -3.0 and -2.0. The models with $k=0.6$ can be ruled out with this restriction.
@@
Figure 12: Distribution of exponential density slopes fitted for the metal-poor MDFs from GCE models exhibiting an $\alpha$/Fe]-rise and from Auriga galaxies measured between $3.0<\mathrm{[Fe/H]}<-2.0$. The median slope for each group is marked with a vertical dashed line with its value annotated. The most common slopes among our GCE models with $\alpha$/Fe]-rise agree with those produced in Auriga. The $\alpha$/Fe]-rise is key to establishing the scenario of the proto-Galaxy and the observed MDF help us further narrow down.
4.2 Robust MDF Slope Across Catalogs and Selections
Figure 5 shows the MDFs over $3<\mathrm{[Fe/H]}<-1.8$ on a logarithmic scale for the Gaia XP catalogs listed in Sec. 2. Using log density highlights the near-exponential falloff of the metal-poor tail, as in Fig. 4 of 2022ApJ...941...45R for $2<\mathrm{[Fe/H]}<-1$. For visual guidance, dashed lines indicate three reference slopes ($k=0.6,1.0,1.4$) that bracket the range we observe across catalogs in this interval. We excluded 2024ApJS..272....2L, whose metallicities are truncated at $\mathrm{[Fe/H]}=-2.0$, producing an artificial pile-up at the boundary.
Because 2024A&A...691A..98K and 2025A&A...695A..75Y are anchored to APOGEE training labels, their MDFs decline rapidly below $\mathrm{[Fe/H]}\approx-2$, where APOGEE offers few measurements. In addition, 2025A&A...695A..75Y disable their APOGEE-based NN flux correction for stars with $\mathrm{[Fe/H]}\leq-2.5$ and fit FERRE directly, which they release as a separate metal-poor subset. This treatment imprints a discontinuity near $\mathrm{[Fe/H]}=-2.5$, but the MDF tail slope remains consistent. 2025ApJS..279....7Y train against metal-poor reference sets (PASTEL, SAGA) with a cost-sensitive objective to enhance recovery of VMP giants, similarly flattening their tail. Finally, 2024A&A...692A.115M-as part of the Pristine effort targeting $\mathrm{[Fe/H]}<-1$-deliver a comparatively larger fraction of metal-poor stars from synthetic CaHK, which also yields a flat density decline in their MDF.
The similarly flat MDF tail in the 2025ApJ...980...90H catalog is best explained by two design choices that boost recovery of VMP halo stars: (1) the training set was deliberately augmented with the VMP sample of 2022ApJ...931..147L, extending label coverage well below $\mathrm{[M/H]}\approx-2.5$ and increasing counts at the low-metallicity end; and (2) the recommended catalog is restricted to low-extinction sight lines, E(BV)<0.1E(B-V)<0.1, which preferentially probe high-latitude halo volume and improves completeness for distant halo giants. However, the same VMP set was also used for training by 2023MNRAS.524.1855Z and 2025MNRAS.537.1984A, yet both deliver noticeably steeper metal-poor tails. This suggests that simply augmenting training set with VMP stars does not necessarily determine the slope; sample selection and quality cuts play a more important role.
2023ApJS..267....8A and 2023MNRAS.524.1855Z appear to under-represent the most metal-poor regime: their MDFs show the steepest slope for $\mathrm{[Fe/H]}<-2.0$ with $k>1.4$, then flatten toward $k\approx 1.0$ as metallicity increases. 2023ApJS..267....8A selected sources in their vetted RGB sample to avoid spurious metal-poor stars, which could preferentially remove some VMP stars. 2023MNRAS.524.1855Z did not explicitly include VMP stars in their training set and their basic reliability cuts tend to cull low-SNR sources, disproportionately trimming the VMP tail. Overall, the metal-poor tails have logarithmic slopes centered near $k\approx 1$, and are well described by a single exponential over our metallicity window. We observe no bumps or peaks within $3<\mathrm{[Fe/H]}<-1.8$ after applying recommended quality cuts from the authors. Minor departures from linearity in log-density appear only at the boundaries-flattening in VMP-boosted catalogs and steepening where training or completeness drops below $\mathrm{[Fe/H]}\approx-2.0$-and are consistent with known selection and pipeline effects.
Figure 6 shows MDFs over $3<\mathrm{[Fe/H]}<-1.8$ split by Galactocentric radius $R$ in (a) and by observed absolute magnitude $G_{\mathrm{obs}}$ in (b), with the same reference slopes from Fig.5. Because the $y$-axis is logarithmic, the curves tend to meet near $\mathrm{[Fe/H]}\simeq-1.8$ at density $\sim 10^{0}$. A quick way to read an individual slope is to note the intercept at $\mathrm{[Fe/H]}=-3$: for example, an intercept of $10^{-1}$ implies $k\approx 1.0/1.5\simeq 0.67$, while $10^{-2}$ implies $k\simeq 1.33$. Most catalogs exhibit nearly identical slopes across the various $R$ and $G_{\mathrm{obs}}$ bins, with only modest deviations.
In Fig.6 (b), several samples contain fewer VMP stars for $4<G_{\mathrm{obs}}<8$. As one moves to fainter $G_{\mathrm{obs}}$, several catalogs approach their effective luminosity limits because their catalogs contain mostly giants, while the rest contain fewer stars. The MDF by 2023MNRAS.524.1855Z loses VMP stars more rapidly at $G_{\mathrm{obs}}\gtrsim 4$, in line with their basic reliability cut, which preferentially culls faint sources with low S/N and high uncertainty. Meanwhile, catalogs from 2025ApJ...980...90H, 2025ApJS..279....7Y, and 2024A&A...692A.115M show essentially constant slopes across $G_{\mathrm{obs}}$ values. The first catalog used VMP stars to supplement their training set and their quality cut relies primarily on a CMD filter which preserves MSTO and MS stars. The other two catalogs benefit from synthetic spectra in their pipeline which mitigates the lack of VMP dwarfs in their training data.
4.3 MDF Slope Tells Early Assembly Tale
We used the GCE model outlined in Sec. 3 from our first paper to explore the physical conditions behind different slopes of the metal-poor tail of MDF. The GCE models presented here are pre-selected in Paper I based on the following criteria: 1) [Fe/H] should reach $1.3\pm 0.05$ dex after one Gyr; 2) [Mg/Fe] should reach $0.26\pm 0.05$ dex after one Gyr; 3) [Mg/Fe] increases above 0.3 dex with a minimum increase of 0.05 dex while [Fe/H] surpasses -0.9 dex in the next Gyr.
@@ 4.3 MDF Slope Tells Early Assembly Tale
Figure 7 shows the median MDFs of the subsets that best match the three reference slopes from Fig. 5. The shaded bands indicate the inter-model spread, and the count of contributing models is annotated beside each $k$ label. Except for $k=1.0$, a single exponential profile is an imperfect description: for $k=0.6$ the MDF steepens toward the metal-rich side once $\mathrm{[Fe/H]}\gtrsim-2.0$, whereas for $k=1.4$ it flattens at the metal-poor end below $\mathrm{[Fe/H]}\lesssim-2.4$. We fitted the MDFs piecewise with sliding windows of 0.5 dex over $3.0<\mathrm{[Fe/H]}<-1.5$. We then retained the windows whose slopes agreed and showed small residuals in log-density, and took their common value as the slope of the metal-poor tail of each MDF. The main limitation is imposed by our GCE setup, specifically the assumption of a constant inflow rate, which we will discuss more in the following. In the resulting model ensemble, the $k=1.0$ category is by far the most populated, with substantially fewer cases in the other two categories.
@@
Figure 10 shows the cumulative stellar mass as a function of metallicity for the three slope classes in our GCE models (Fig. 7). We construct this diagnostic because ages for the oldest stars are notoriously uncertain, whereas $\rm[Fe/H]$} is directly observable at scale from Gaia XP. As a reference point, we mark the estimated stellar mass of the in-situ proto-Galaxy (the Aurora population) from 2024arXiv241022250K, who found a mass between $1.07\times 10^{7}$ and $2.11\times 10^{7}\,\mathrm{M_{\odot}}$ for stars in the range $3.0<\mathrm{[Fe/H]}<-2.0$. This mass was derived from modeling the spatial density distribution of metal-poor red giant branch stars from Gaia, accounting for the survey’s selection function.
@@
The curve shapes encode the same information highlighted in Fig. 9: in the burst-quenched case ($k=0.6$), an early starburst builds $\sim 10^{7}\,{\rm M_{\odot}}$ of stars rapidly, but the large initial gas reservoir keeps the initial metallicity low, producing a fast early rise that soon plateaus. In contrast, the slow-assembly case ($k=1.4$) begins with a milder burst into a less massive reservoir, reaching a higher initial metallicity and then growing steadily under sustained low-level inflow and higher SFE; the most common slopes in both GCE models and Auriga galaxies ($k\simeq 1.2$) lie between these limits.
4.4 Quiet Dawn Leads to Unified Slope
@@
Auriga adopts Arepo’s Galilean-invariant, Voronoi moving-mesh hydrodynamics–crucial for accurately capturing shocks, contact discontinuities, and rotating disks–and builds on a comprehensive galaxy-formation model (cooling, star formation, stellar/AGN feedback, metal enrichment) inherited from the Illustris framework (2010MNRAS.401..791S; 2013MNRAS.436.3031V). The simulations self-consistently amplify magnetic fields and produce disk galaxies whose global properties (e.g., Hi distributions, kinematics, scaling relations) compare well with observations, providing a physically grounded laboratory for disk formation and chemo-dynamical studies (2017MNRAS.466.3859M; 2017MNRAS.469.3185P). A recent overview details the augmented suite and the first public data release–raw snapshots, group catalogs, merger trees, and analysis tools [^1] https://wwwmpa.mpa-garching.mpg.de/auriga/data.html–enabling community access to the methodology and outputs (10.1093/mnras/stae1598).
Figure 11 shows the time evolution of global properties for the 30 Auriga galaxies in the original release, in the same style as Fig. 9. Panels display: (a) total gas mass; (b) SFR; (c) cumulative stellar mass; (d) mass-weighted stellar [Fe/H]; and (e) mass-weighted stellar [Mg/Fe]. The commonly used Milky Way analogues-Au-17, Au-18, Au-23, and Au-26-are highlighted in orange; all other galaxies are shown in gray. Au-17 and Au-18 best reproduce the Milky Way bulge/bar chemo-kinematics and exhibit quiescent merger histories since $z\!\sim\!3.5$ (2020MNRAS.494.5936F). Au-18 also experiences an early Gaia-Sausage/Enceladus-like merger and has halo properties closest to the Milky Way (2024MNRAS.531.1520M). Au-17 also undergoes a similar merger, but its properties are less consistent with those of GSE (2019MNRAS.484.4471F; 2025arXiv250909576O).
Au-18, Au-23, and Au-26 possess Milky Way-like bars and have been used to time the disk/bar formation in our Galaxy (2024MNRAS.535.2873Z). To align the simulation with the stellar disk, the z-axis is defined by the net angular momentum vector of all stars within 0.1 times the virial radius ($0.1\times r_{200c}$). We restricted the analysis to stars with $R<10$ kpc and $|\mathrm{z}|<3$ kpc to match the spatial footprint of Gaia XP catalogs. Comparing Figs. 9 and 11, most Auriga systems are consistent with the k=1.0k\!=\!1.0 GCE category, characterized by negligible initial gas reservoirs and minimal initial star formation. For each galaxy, we constructed MDFs from the masses and [Fe/H] of surviving star particles at $z\!\sim\!0$ and estimated the slope of the metal-poor tail over $3<\mathrm{[Fe/H]}<-1.8$.
@@
5 discussion
5.1 A Unified Proto-Galactic Population
We interpret the proto-Galaxy as the composite stellar population that formed up to the transition from dispersion-dominated assembly to the emergence of a rotation-supported disk. Across Gaia XP metallicity catalogs, the metal-poor MDF tails ($3<\mathrm{[Fe/H]}<-2.0$) follow nearly single-slope exponentials ($k\approx 1$), consistent with a chemically well-mixed system whose early star formation was regulated by gas supply rather than distinct subcomponents.
Over the metallicity interval, the metal-poor tails of the MDFs from various Gaia XP catalogs follow an approximately single-slope exponential density with logarithmic slopes centered near $k\approx 1$. After applying each catalog’s recommended quality cuts, no secondary peaks or shoulders are detected within this range. Modest departures occur only at the edges and are attributable to known pipeline and selection boundaries: APOGEE-anchored products steepen below $\mathrm{[Fe/H]}=-2$ because of scarce training labels, whereas VMP-boosted designs (e.g., Ca HK-based or explicitly VMP-augmented training) flatten the extreme metal-poor tail.
@@ 5.1 A Unified Proto-Galactic Population
Both observations and simulations show that such old, metal-poor stars arise from a mixture of origins: an in-situ component heated by early mergers and accreted debris from one or two major progenitors (2010ApJ...708.1398T; 2018MNRAS.480..652E; 2024MNRAS.527.9810H). However, disentangling these channels is difficult even in simulations with perfect birth tags-kinematics and chemistry overlap strongly below $\mathrm{[Fe/H]}<-1$, with mean abundance offsets $\lesssim 0.3$ dex and broad MDF overlap in both prograde and retrograde regions (2021MNRAS.503.5846R; 2023A&A...677A..91K). Observational separations based on integrals of motion recover only a small subset of accreted structures (2022ApJ...937...14B), and recent FIRE-2 analyses show significant kinematic and chemical blending during the proto-Galactic phase (2024MNRAS.527.9810H).
@@
5.2 Addressing Caveats on MDFs
@@
A related consideration concerns the implicit selection function of stars from the GCE models and its connection to the observed CMD coverage of Gaia XP catalogs. The MDF represents stars that have survived to the present day, effectively sampling the least massive, longest-lived stars. In principle, differences among catalogs in their CMD coverage–for example, whether they focus on main-sequence dwarfs or red-giant branches–could bias the inferred MDF if stars of different masses contributed differently to the metal-poor tail. However, inspection of the GCE model outputs shows that nearly all stars surviving to $z=0$ originate from one or a few of the lowest-mass bins in the adopted IMF, since higher-mass stars have long evolved and recycled their metals into the ISM.
@@
Consequently, the shape of the MDF is governed almost entirely by the cumulative star formation and enrichment history rather than by the present-day stellar mass distribution. Because all XP-based catalogs ultimately probe these same long-lived low-mass populations, their differing CMD footprints do not systematically affect the inferred slope $k$. In this regime, the MDF can be interpreted as a faithful tracer of the proto-Galaxy’s early gas-regulation history, largely insensitive to catalog-specific evolutionary or photometric selections. Nevertheless, we implemented different selection functions based on Gaia magnitudes of the stellar component in our models and found that the resulting MDFs are identical. This agrees with panels (b) in Fig.6 in which MDF slopes remain consistent across different magnitude bins for catalogs with sufficient stars.
Similarly, the distinct spatial footprints of the Gaia XP catalogs could in principle influence the MDF. However, the MDF slopes remain consistent across Galactocentric radius in Fig. 6, indicating that distance selection has little effect on the slope measurement. Within the one-zone framework adopted here, the proto-Galaxy is approximated as a chemically well-mixed reservoir during its earliest phase, when short dynamical and mixing timescales erased spatial abundance gradients. In this regime, local distance variations trace the same global chemical evolution history, and therefore do not modify the intrinsic shape of the metal-poor MDF.
A further caveat involves the rotation velocities of stars within our metallicity range. Kinematic diagnostics confirm that stars with $\mathrm{[Fe/H]}\lesssim-1.5$ exhibit near-zero median rotation ($v_{\phi}$), consistent with expectations for a proto-Galactic population. However, several Gaia XP catalogs display a secondary concentration of stars with $v_{\phi}>200~\mathrm{km\,s^{-1}}$ at $\mathrm{[Fe/H]}>-1.5$ (Fig. 2). Stars following the isochrone of old, metal-poor populations on the CMD show low $|v_{\phi}|$, whereas those with high $v_{\phi}$ deviate from this locus and occupy CMD regions sparsely represented in the training sets of many XP-based pipelines (Fig. 4). These fast rotators are best interpreted as disk contaminants whose metallicities have been underestimated, rather than genuine very metal-poor stars with large angular momentum. Removing such interlopers requires high-quality radial velocities, but only $\sim$5–10% of XP sources have Gaia RVS measurements, limiting a clean kinematic excision. The small fraction of potential contaminants and their confinement to the metal-richer boundary, however, suggest that they have minimal impact on the MDF slope within $3<\mathrm{[Fe/H]}<-2.0$.
5.3 Comparison with Previous Halo MDF Studies

### v2

@@ IV.3 MDF Slope Tells Early Assembly Tale
Figure 8: Median MDFs from one-zone GCE models grouped by nearest reference MDF class corresponding to nominal slope $k=0.6,1.0,1.4$. The number of models within each slope category is annotated. Lines and shaded regions show the median and 16th–84th percentile ranges, respectively.
@@
Figure 11: The cumulative stellar mass assembly history as a function of metallicity ([Fe/H]) for three slope classes in our GCE models. Lines and shaded regions show the median and 16th–84th percentile ranges, respectively. Kurbatov et al. (2024) estimated the mass of the in situ proto-Galaxy to be between $1.07\times 10^{7}$ and $2.11\times 10^{7}\,\mathrm{M_{\odot}}$ between [Fe/H] = -3.0 and -2.0. The models with $k=0.6$ can be ruled out with this restriction.
@@
Figure 13: Distribution of exponential density slopes fitted for the metal-poor MDFs from GCE models exhibiting an $\alpha$/Fe]-rise and from Auriga galaxies measured between $3.0<\mathrm{[Fe/H]}<-2.0$. The histograms are computed using a uniform bin width of 0.025 in slope. The median slope for each group is marked with a vertical dashed line with its value annotated. The most common slopes among our GCE models with $\alpha$/Fe]-rise agree with those produced in Auriga. The $\alpha$/Fe]-rise is key to establishing the scenario of the proto-Galaxy and the observed MDF help us further narrow down.
@@
Figure 8 shows the median MDFs of the subsets of GCE models that best match the three reference slopes from Fig. 5. The shaded bands indicate the inter-model spread, and the count of contributing models is annotated beside each $k$ label. Except for $k=1.0$, a single exponential profile is an imperfect description: for $k=0.6$ the MDF steepens toward the metal-rich side once $\mathrm{[Fe/H]}\gtrsim-2.0$, whereas for $k=1.4$ it flattens at the metal-poor end below $\mathrm{[Fe/H]}\lesssim-2.4$. We fitted the MDFs piecewise with sliding windows of 0.5 dex over $3.0<\mathrm{[Fe/H]}<-1.5$. We then retained the windows whose slopes agreed and showed small residuals in log-density, and took their common value as the slope of the metal-poor tail of each MDF. The main limitation is imposed by our GCE setup, specifically the assumption of a constant inflow rate, which we will discuss more in the following. In the resulting model ensemble, the $k=1.0$ category is by far the most populated, with substantially fewer cases in the other two categories.
@@
Figure 11 shows the cumulative stellar mass as a function of metallicity for the three slope classes in our GCE models (Fig. 8). We construct this diagnostic because ages for the oldest stars are notoriously uncertain, whereas $\rm[Fe/H]$} is directly observable at scale from Gaia XP. As a reference point, we mark the estimated stellar mass of the in-situ proto-Galaxy (the Aurora population) from Kurbatov et al. (2024), who found a mass between $1.07\times 10^{7}$ and $2.11\times 10^{7}\,\mathrm{M_{\odot}}$ for stars in the range $3.0<\mathrm{[Fe/H]}<-2.0$. This mass was derived from modeling the spatial density distribution of metal-poor red giant branch stars from Gaia, accounting for the survey’s selection function.
@@
IV.4 Quiet Dawn Leads to Unified Slope
@@
Auriga adopts Arepo’s Galilean-invariant, Voronoi moving-mesh hydrodynamics–crucial for accurately capturing shocks, contact discontinuities, and rotating disks–and builds on a comprehensive galaxy-formation model (cooling, star formation, stellar/AGN feedback, metal enrichment) inherited from the Illustris framework (Springel, 2010; Vogelsberger et al., 2013). The simulations self-consistently amplify magnetic fields and produce disk galaxies whose global properties (e.g., Hi distributions, kinematics, scaling relations) compare well with observations, providing a physically grounded laboratory for disk formation and chemo-dynamical studies (Marinacci et al., 2017; Pakmor et al., 2017). A recent overview details the augmented suite and the first public data release–raw snapshots, group catalogs, merger trees, and analysis tools [^3] https://wwwmpa.mpa-garching.mpg.de/auriga/data.html–enabling community access to the methodology and outputs (Grand et al., 2024).
Figure 12 shows the time evolution of global properties for the 30 Auriga galaxies in the original release, in the same style as Fig. 10. Panels display: (a) total gas mass; (b) SFR; (c) cumulative stellar mass; (d) mass-weighted stellar [Fe/H]; and (e) mass-weighted stellar [Mg/Fe]. The commonly used Milky Way analogues-Au-17, Au-18, Au-23, and Au-26-are highlighted in orange; all other galaxies are shown in gray. Au-17 and Au-18 best reproduce the Milky Way bulge/bar chemo-kinematics and exhibit quiescent merger histories since $z\!\sim\!3.5$ (Fragkoudi et al., 2020). Au-18 also experiences an early Gaia-Sausage/Enceladus-like merger and has halo properties closest to the Milky Way (Merrow et al., 2024). Au-17 also undergoes a similar merger, but its properties are less consistent with those of GSE (Fattahi et al., 2019; Orkney & Laporte, 2025).
Au-18, Au-23, and Au-26 possess Milky Way-like bars and have been used to time the disk/bar formation in our Galaxy (Zhang et al., 2024). To align the simulation with the stellar disk, the z-axis is defined by the net angular momentum vector of all stars within 0.1 times the virial radius ($0.1\times r_{200c}$). We restricted the analysis to stars with $R<10$ kpc and $|\mathrm{z}|<3$ kpc to match the spatial footprint of Gaia XP catalogs. Comparing Figs. 10 and 12, most Auriga systems are consistent with the k=1.0k\!=\!1.0 GCE category, characterized by negligible initial gas reservoirs and minimal initial star formation. For each galaxy, we constructed MDFs from the masses and [Fe/H] of surviving star particles at $z\!\sim\!0$ and estimated the slope of the metal-poor tail over $3<\mathrm{[Fe/H]}<-1.8$.
@@
V discussion
V.1 Selection function considerations
The Gaia XP metallicity catalogs considered in this work do not share a single selection function. In addition to the native Gaia XP magnitude limit, each catalog has its own effective selection through the choice of training set, inference domain, and recommended quality cuts. Throughout this work, we therefore do not impose additional cuts of our own beyond the published recommended quality cuts; our goal is to compare the public catalogs within their validated regimes rather than construct a new common sample with a different and difficult-to-model selection function.
A related issue is evolutionary state. Because metal-poor giants are intrinsically brighter than dwarfs, giant-dominated catalogs probe larger distances and a different mixture of Galactic components than catalogs that include substantial numbers of dwarfs and subgiants. We do not apply a uniform giant/dwarf cut across catalogs, since several public releases are already restricted to vetted giant samples whereas others are intended to operate over broader CMD regions. Instead, we assess the effect of this heterogeneity empirically. As shown above, the logarithmic MDF slope remains broadly consistent across catalogs with different CMD footprints and also across bins of observed absolute magnitude and Galactocentric radius (Figs. 5,  7, and 6) indicating that luminosity-class and volume effects primarily alter the normalization and sampled volume rather than strongly changing the fitted MDF slope over the adopted metallicity interval.
Disk contamination is another concern because the Gaia XP samples span both low- and high-$|\mathrm{z}|$ regions. We do not apply a global kinematic cut to the catalogs because only a minority of stars ($\sim$5–10%) have Gaia DR3 RVS radial velocities; instead, we use this subset only as a diagnostic. The available kinematics show that stars with $\mathrm{[Fe/H]}<-2.0$ are predominantly slow-rotating, consistent with a proto-Galactic population, whereas the clearest fast-rotating contaminants become common closer to the metal-richer boundary, especially above $\mathrm{[Fe/H]}\gtrsim-1.5$ (Fig. 2). These stars often deviate from the old, metal-poor isochrone locus in the CMD and occupy regions that are sparsely represented in the training sets of several XP-based pipelines (Fig. 4), indicating that they are more likely metal-rich disk contaminants with underestimated metallicities than genuine very metal-poor stars with large angular momentum. Catalogs with larger low-$|z|$ samples are therefore more susceptible to this effect (see the last panel of Fig. 2 and Fig. 3). To minimize such contamination, we restrict the MDF-tail analysis to $\mathrm{[Fe/H]}<-2.0$ and treat the RVS-based kinematic distributions only as a validation check rather than as a selection criterion. The confinement of the clearest contaminants to the metal-richer boundary suggests that their impact on the fitted MDF slope over $3<\mathrm{[Fe/H]}<-2.0$ is small.
V.2 A Unified Proto-Galactic Population
We interpret the proto-Galaxy as the composite stellar population that formed up to the transition from dispersion-dominated assembly to the emergence of a rotation-supported disk. Across the Gaia XP metallicity catalogs, the metal-poor MDF tails over $3<\mathrm{[Fe/H]}<-2.0$ are generally well described by an approximately single-slope exponential with logarithmic slopes centered near $k\approx 1$. After applying each catalog’s recommended quality cuts, we find no compelling evidence for secondary peaks or shoulders within this interval. Residual departures appear mainly near the boundaries and are consistent with known pipeline and selection effects: APOGEE-anchored products tend to steepen below $\mathrm{[Fe/H]}\approx-2$ because of sparse metal-poor training labels, whereas VMP-optimized or CaHK-based designs can flatten the extreme metal-poor end. These effects likely account for much of the inter-catalog variation in the fitted slopes, but they do not qualitatively change the main result that the metal-poor tail is broadly smooth and close to exponential over the adopted metallicity range.
@@ V.2 A Unified Proto-Galactic Population
Both observations and simulations show that such old, metal-poor stars arise from a mixture of origins: an in-situ component heated by early mergers and accreted debris from one or two major progenitors (Tumlinson, 2010; El-Badry et al., 2018; Horta et al., 2024). However, disentangling these channels is difficult even in simulations with perfect birth tags-kinematics and chemistry overlap strongly below $\mathrm{[Fe/H]}<-1$, with mean abundance offsets $\lesssim 0.3$ dex and broad MDF overlap in both prograde and retrograde regions (Renaud et al., 2021; Khoperskov et al., 2023). Observational separations based on integrals of motion recover only a small subset of accreted structures (Brauer et al., 2022), and recent FIRE-2 analyses show significant kinematic and chemical blending during the proto-Galactic phase (Horta et al., 2024).
@@
V.3 Addressing Caveats on MDFs
@@
A related consideration concerns the implicit selection function of stars from the GCE models and its connection to the observed CMD coverage of Gaia XP catalogs. The MDF represents stars that have survived to the present day, effectively sampling the least massive, longest-lived stars. In principle, differences among catalogs in their CMD coverage–for example, whether they focus on main-sequence dwarfs or red-giant branches–could bias the inferred MDF if stars of different masses contributed differently to the metal-poor tail. However, inspection of the GCE model outputs shows that nearly all stars surviving to $z=0$ originate from one or a few of the lowest-mass bins in the adopted IMF ($\mathrm{M}_{*}<0.7\mathrm{M}_{\odot}$), since higher-mass stars have long evolved and recycled their metals into the ISM.
@@
Consequently, the shape of the MDF is governed almost entirely by the cumulative star formation and enrichment history rather than by the present-day stellar mass distribution. Because all XP-based catalogs ultimately probe these same long-lived low-mass populations, their differing CMD footprints do not systematically affect the inferred slope $k$. In this regime, the MDF can be interpreted as a faithful tracer of the proto-Galaxy’s early gas-regulation history, largely insensitive to catalog-specific evolutionary or photometric selections. Nevertheless, we implemented different selection functions based on Gaia magnitudes of the stellar component in our models and found that the resulting MDFs are identical. This agrees with Fig. 7 in which MDF slopes remain consistent across different parallax-corrected magnitude bins for catalogs with sufficient stars.
V.4 Comparison with Previous Halo MDF Studies